OpenStudy (anonymous):

Two forces with magnitudes of 200 and 100 pounds act on an object at angles of 60° and 170° respectively. Find the direction and magnitude of these forces. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

u=200[cos(60), sin(170)] =u=-81.028 I've done this much so far but I have no idea if it's correct!

OpenStudy (anonymous):

draw a line of 10cm at an angle 60 degree to x axis and another 5 cm at an angle 170 degrees. Add the tip of 5cm to the tip os 10 cm. Measure the length and the angle made by it wrt x axis. Multiply by 20 to it. This will give the magnitude and the angle will be that of the resultant.

OpenStudy (anonymous):

@amistre64 please help me :(

OpenStudy (anonymous):

@Mertsj @lalaly I've been stuck on this forever!

OpenStudy (anonymous):

@RadEn help

OpenStudy (mertsj):

u=200 at 60 degrees = 200 cos 60i + 200 sin 60j=100i + 173.2j v=100 at 170 degrees = 100 cos 170i + 100 sin 170j=-98.4i+17.4j u+v=1.6i+190.6j

OpenStudy (mertsj):



OpenStudy (anonymous):

Lol and I thought I was close.... but I'm still confused on how to find the direction?

OpenStudy (mertsj):

tanx=190.6/1.6=119.1 x=89.5 degrees sqrt[(1.6)^2+(190.6)^2]=190.6 So the new vector is magnitude 190.6 at a direction of 89.6 degrees. I did not check my arithmetic. You can do that.

OpenStudy (anonymous):

Oh my this is confusing but thanks a ton @Mertsj ! (:

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Can you help me with one more problem? This time it's easier! @Mertsj

OpenStudy (mertsj):

What is the problem?

OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7> Idk why but it's giving me so much trouble!